# Controlling the Flow: Stability and Convergence for Stochastic Gradient Descent with Decaying Regularization

### v1

@@ A.2 A toy example
with unique minimum-norm solution $x_{\ast}=(1/2,1/2)$, see the plot in Section 1.2. Note that there exist infinitely many global minima of $f$. Incorporating Tikhonov regularization results in
@@
Implementation details: We have implemented both vanilla SGD and reg-SGD by hand and initialized both algorithms with same initial state $X_{0}\sim\mathcal{N}(0,1)$ and perturbed the exact gradient $\nabla f$ in each iteration by independent noise following a multivariate normal distribution with zero mean and covariance $0.1^{2}\cdot\mathrm{Id}$. For SGD we chose the step-size schedule $\alpha_{k}=0.1k^{-1/2}$, $k\in\mathbb{N}$. For reg-SGD we chose $\alpha_{k}=0.1k^{-q}$ and regularization $\lambda_{k}=k^{-p}$, where $p=\frac{1}{4\xi+3}$, $q=(1+p)/2$ when considering the $L^{2}$ convergence rates and $p=(6\xi+3)^{-1}$, $q=2/3$ when considering the almost sure convergence rates see Corollary 2.6 and Corollary 2.7.
@@
The plots of Figures 6 and 7 illustrate that reg-SGD converges to the minimum-norm solution both in $L^{2}$ (Figure 6) and almost surely (Figure 7), as indicated by the vanishing squared error. In contrast, SGD does not converge to the minimum-norm solution, although it achieves convergence in the expected (Figure 6) and pathwise optimality gap (Figure 7). This highlights the regularization effect of reg-SGD in guiding the iterates toward the unique minimum-norm solution as also indicated in Figure 9.
@@
Figure 9: Convergence rate for $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]$ in the situation of Corollary 2.6 (left) and almost sure convergence for $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ in the situation of Corollary 2.7 in the considered setting of Section A.2 with $\xi=1$. Furthermore, we display the choices $(p,q)\in\{(0.111,0.667),(0,0.667),(0.67,0.5),(0.111,0.29)\}$ which are simulated and displayed in Figure 8.
Finally, we compare different choices of $(p,q)$ for reg-SGD. In particular, we run reg-SGD with $\alpha_{k}=0.2qk^{-1/2}$ and $\lambda_{k}=k^{-p}$ for the choices
@@
Moreover, we increase the noise covariance to $\mathcal{N}(0,\mathrm{Id})$. The expected convergence behavior is shown in Figure 9 while the resulting errors are displayed in Figure 8. As expected, we do not observe convergence for the choices $(0,0.667)$ and $(0.67,0.5)$ as the regularization is not turned off, respectively turned off too fast. We observe convergence both a.s. and in $L^{2}$ when choosing $(0.111,0.667)$ as suggested by our theory. In contrast, when choosing $(0.111,0.29)$ our theoretical results suggest that the step-size decay is too slow, which we observe in a high variance of the deviation to the minimum-norm solution.
@@
We implemented both SGD and reg-SGD by partitioning the forward operator $A\in\mathbb{R}^{K\times d}$ into rows $A_{i}\in\mathbb{R}^{1\times d}$, $i=1,\dots,K$. Hence, $A_{i}x$ corresponds to the discretized ODE solution at location $s_{i}$. At each iteration, a batch of $16$ locations $(s_{i_{1}},\dots,s_{i_{16}})$ are sampled uniformly at random, and the gradient of $f$ is approximated by $\nabla f(x)\approx\frac{1}{16}\sum_{j=1}^{16}A_{i_{j}}^{\top}(A_{i_{j}}x-y_{i_{j}})\in\mathbb{R}^{d}$ and additionally perturbed by independent noise following a multivariate normal distribution with zero mean and covariance $0.001^{2}\cdot{\mathrm{Id}}$. For SGD we chose the learning rate schedule $\alpha_{k}=100k^{-1/2}$. For reg-SGD we chose $\alpha_{k}=100k^{-2/3}$ and regularization $\lambda_{k}=0.001k^{-1/3}$. We initialize both algorithms for each repetition at zero.
@@
In Figure 10, we compare the expected and pathwise optimality gap (left) as well as the $L^{2}$ and pathwise error to the minimum-norm solution (right). While SGD shows fast convergence in terms of the optimality gap, it again fails to converge to the minimum-norm solution. In contrast, reg-SGD slows down the convergence in terms of the optimality gap, but safely reconstructs the minimum-norm solution. In Figure 11 (left), we plot the reconstruction of the unknown right-hand side $x^{\dagger}$ resulting from the minimum-norm solution $x_{\ast}=A^{\dagger}y$, and from the last iterates of SGD and reg-SGD. Moreover, in Figure 11 (right) we plot the corresponding ODE solutions when solving (10) with the estimated right-hand side.
@@ Appendix B Auxiliary results
Suppose that $f$ satisfies 1.1, then the following statements hold true:
@@
Under 1.1 the following estimates are satisfied for all $x\in\mathcal{X}$ and $\lambda\geq 0$:
@@
Lemma B.5 (See e.g. Lemma A.3. in [54]).
Suppose that $(z_{k})_{k\in\mathbb{N}_{0}}$ is a positive sequence such that $z_{k}\leq\left(1-a_{k}\right)z_{k-1}+c_{k}$ for all $k\in\mathbb{N}$, where $(a_{k})_{k\in\mathbb{N}},(c_{k})_{k\in\mathbb{N}}$ are positive sequences satisfying
$\sum_{k=1}^{\infty}a_{k}=\infty\quad\text{ and }\quad\sum_{k=1}^{\infty}c_{k}<\infty.$
Then $\lim_{k\rightarrow\infty}z_{k}=0$.
@@
Let $C\geq 0$ be a constant such that for all $k\in\mathbb{N}$ one has $Y_{k}\geq-C$, almost surely. Set $(\tilde{Y}_{k})_{k\in\mathbb{N}}=(Y_{k}+C)_{k\in\mathbb{N}}$ and note that (12) still holds when replacing $(Y_{k})_{k\in\mathbb{N}}$ by $(\tilde{Y}_{k})_{k\in\mathbb{N}}$. Thus, the statement follows from Lemma B.6 for the choice $\gamma_{k}\equiv 0$, $X_{k}\equiv 0$ and $(Z_{k})_{k\in\mathbb{N}}=(z_{k})_{k\in\mathbb{N}}$. ∎
Corollary B.8 (See e.g. Lemma A.2 in [54]).
Let $(\mathcal{F}_{k})_{k\in\mathbb{N}}$ be a filtration and $(Y_{k})_{k\in\mathbb{N}}$, $(A_{k})_{k\in\mathbb{N}}$ and $(B_{k})_{k\in\mathbb{N}}$ be non-negative and adapted processes satisfying almost surely that
$\sum_{k=1}^{\infty}A_{k}=\infty\quad\text{ and }\quad\sum_{k=1}^{\infty}B_{k}<\infty.$
@@
Appendix C Proofs of the main results
@@
Given Lemma B.3, this process $(E_{k})_{k\in\mathbb{N}_{0}}$ serves as Lyapunov function for computing the convergence rates stated in Section 2.
Proposition C.1.
Suppose that 1.1 and 1.2 are fulfilled and and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with predictable (random) step-sizes and regularization parameters that are uniformly bounded from above and such that $(\lambda_{k})_{k\in\mathbb{N}}$ is almost surely decreasing. For $k\in\mathbb{N}$ denote by $\mathbb{A}_{k}=\{\alpha_{k}\leq\frac{2}{L+\lambda_{k}}\}\in\mathcal{F}_{k-1}$. Then, for all $k\in\mathbb{N}$,
$\mathbb{E}[1\mathrm{l}_{\mathbb{A}_{k}}E_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Big{(}1-2\lambda_{k}\alpha_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big{)}1\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}$
@@
Using 1.2, the property $f\leq f_{\lambda_{k}}$, and the descent condition (11) applied to the $(L+\lambda_{k})$-smooth function $f_{\lambda_{k}}$, yields, for $k\in\mathbb{N}$,
$\mathbb{E}[1\mathrm{l}_{\mathbb{A}_{k}}f_{\lambda_{k}}(X_{k})\mid\mathcal{F}_{k-1}]$ $\leq 1\mathrm{l}_{\mathbb{A}_{k}}\Big{(}f_{\lambda_{k}}(X_{k-1})-\alpha_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}\|\nabla f_{\lambda_{k}}(X_{k-1})\|_{\mathcal{X}}^{2}$
@@
$\leq 1\mathrm{l}_{\mathbb{A}_{k}}\Big{(}f_{\lambda_{k}}(X_{k-1})-\alpha_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}\|\nabla f_{\lambda_{k}}(X_{k-1})\|_{\mathcal{X}}^{2}$
@@
$\leq 1\mathrm{l}_{\mathbb{A}_{k}}\Big{(}f_{\lambda_{k}}(X_{k-1})-\alpha_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}\|\nabla f_{\lambda_{k}}(X_{k-1})\|_{\mathcal{X}}^{2}$
@@
where in the last step we also used Lemma B.1 (iii). Since each $f_{\lambda_{k}}$ is $\lambda_{k}$-strongly convex, it satisfies the Polyak-Łojasiewicz inequality
@@
$\mathbb{E}[1\mathrm{l}_{\mathbb{A}_{k}}f_{\lambda_{k}}(X_{k})\mid\mathcal{F}_{k-1}]$ $\leq 1\mathrm{l}_{\mathbb{A}_{k}}\Bigl{(}f_{\lambda_{k}}(X_{k-1})-2\alpha_{k}\lambda_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}(f_{\lambda_{k}}(X_{k-1})-f_{\lambda_{k}}(x_{\lambda_{k}}))$
@@
since $f_{\lambda_{k+1}}(X_{k})-f_{\lambda_{k}}(X_{k})\leq 0$. Combining the previous computations and using Lemma B.1 (ii) yields
$\mathbb{E}[1\mathrm{l}_{\mathbb{A}_{k}}E_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Bigl{(}1-2\lambda_{k}\alpha_{k}\Bigl{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Bigr{)}\Bigr{)}1\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}+f_{\lambda_{k}}(x_{\lambda_{k}})-f_{\lambda_{k+1}}(x_{\lambda_{k+1}})$
$\quad+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}\bigl{(}1\mathrm{l}_{\mathbb{A}_{k}}A(f_{\lambda_{k}}(X_{k-1})-f_{\lambda_{k}}(x_{\lambda_{k}}))+A\frac{\lambda_{k}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+C\bigr{)}$
$\leq\Bigl{(}1-2\lambda_{k}\alpha_{k}\Bigl{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Bigr{)}\Bigr{)}1\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}+\frac{\lambda_{k}-\lambda_{k+1}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}$
$\quad+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}\Bigl{(}1\mathrm{l}_{\mathbb{A}_{k}}AE_{k-1}+A\frac{\lambda_{k}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+C\Bigr{)}.$
With the help of the energy function $(E_{k})_{k\in\mathbb{N}_{0}}$, we can bound the optimality gap of the true objective function, as well as the distance to the unique minimizer of the regularized objective function. For this, we rephrase Lemma B.3 in the notation used in this section.
Lemma C.2.
Suppose that 1.1 is fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4). Then the following estimates are satisfied for all $k\in\mathbb{N}$:
@@
    $\|X_{k}-x_{\lambda_{k}}\|_{\mathcal{X}}^{2}\leq\frac{2E_{k}}{\lambda_{k+1}}$.
C.1 General convergence result
First, we prove the general convergence results. We emphasize that no boundedness assumption is imposed on the reg-SGD scheme. In fact, in the proof of Theorem 2.1 we will show that Assumptions 1.1-1.2 together with (15) imply that $\sup_{k\in\mathbb{N}_{0}}\|X_{k}\|_{\mathcal{X}}^{2}<\infty$ almost surely. The proof uses ideas from Theorem 4.1 in [40] and Lemma 3.1 in [19]. Due to the discretization error, we introduce and analyze a combined Lyapunov function $(\varphi_{k}+E_{k})_{k\in\mathbb{N}}$, where $\varphi_{k}=\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ and $E_{k}$ is defined in (13), in order to prove a descent step.
C.1.1 Proof of Theorem 2.1

### v2

@@ A.2 A toy example
with unique minimum-norm solution $x_{\ast}=(1/2,1/2)$, see the plot in Section˜1.2. Note that there exist infinitely many global minima of $f$. Incorporating Tikhonov regularization results in
@@
Implementation details: We have implemented both vanilla SGD and reg-SGD by hand and initialized both algorithms with same initial state $X_{0}\sim\mathcal{N}(0,1)$ and perturbed the exact gradient $\nabla f$ in each iteration by independent noise following a multivariate normal distribution with zero mean and covariance $0.1^{2}\cdot\mathrm{Id}$. For SGD we chose the step-size schedule $\alpha_{k}=0.1k^{-1/2}$, $k\in\mathbb{N}$. For reg-SGD we chose $\alpha_{k}=0.1k^{-q}$ and regularization $\lambda_{k}=k^{-p}$, where $p=\frac{1}{4\xi+3}$, $q=(1+p)/2$ when considering the $L^{2}$ convergence rates and $p=(6\xi+3)^{-1}$, $q=2/3$ when considering the almost sure convergence rates see Section˜2.3 and Section˜2.3.
@@
The plots of Figures 6 and 7 illustrate that reg-SGD converges to the minimum-norm solution both in $L^{2}$ (Figure˜6) and almost surely (Figure˜7), as indicated by the vanishing squared error. In contrast, SGD does not converge to the minimum-norm solution, although it achieves convergence in the expected (Figure˜6) and pathwise optimality gap (Figure˜7). This highlights the regularization effect of reg-SGD in guiding the iterates toward the unique minimum-norm solution as also indicated in Figure˜1.
Figure 8: Convergence rate for $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]$ in the situation of Section˜2.3 (left) and almost sure convergence for $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ in the situation of Section˜2.3 in the considered setting of Section˜A.2 with $\xi=1$. Furthermore, we display the choices $(p,q)\in\{(0.111,0.667),(0,0.667),(0.67,0.5),(0.111,0.29)\}$ which are simulated and displayed in Figure˜9.
@@
Next, we compare different choices of $(p,q)$ for reg-SGD. In particular, we run reg-SGD with $\alpha_{k}=0.2qk^{-1/2}$ and $\lambda_{k}=k^{-p}$ for the choices
@@
Moreover, we increase the noise covariance to $\mathcal{N}(0,\mathrm{Id})$. The expected convergence behavior is shown in Figure˜8 while the resulting errors are displayed in Figure˜9. As expected, we do not observe convergence for the choices $(0,0.667)$ and $(0.67,0.5)$ as the regularization is not turned off, respectively turned off too fast. We observe convergence both a.s. and in $L^{2}$ when choosing $(0.111,0.667)$ as suggested by our theory. In contrast, when choosing $(0.111,0.29)$ our theoretical results suggest that the step-size decay is too slow, which we observe in a high variance of the deviation to the minimum-norm solution. In the final experiment, we examine the effect of the initial value $\alpha_{1}>0$ in the step-size schedule. For SGD, we set $\alpha_{k}=\alpha_{1}k^{-1/2}$, while for reg-SGD we fix $\lambda_{k}=k^{-0.111}$ and use the step-size schedule $\alpha_{k}=\alpha_{1}k^{-0.667}$. We report both the pathwise optimality gap and the pathwise squared error to the minimum-norm solution for SGD (Figure˜10) and reg-SGD (Figure˜11) under varying initial step sizes $\alpha_{1}\in\{0.01,0.1,1,2\}$.
Figure 10: Left: pathwise optimality gap $f(X_{k})-f(x_{\ast})$. Right: pathwise squared error to the minimum-norm solution $\|X_{k}-x_{\ast}\|^{2}$. Each curve represents one of $10$ independent runs of SGD, each of length $N=10^{7}$. The shaded lines depict individual runs of SGD. The solid lines show the average errors for SGD. The different colors correspond to various choices of $\alpha_{1}\in\{0.01,0.1,1,2\}$.
Figure 11: Left: pathwise optimality gap $f(X_{k})-f(x_{\ast})$. Right: pathwise squared error to the minimum-norm solution $\|X_{k}-x_{\ast}\|^{2}$. Each curve represents one of $10$ independent runs of reg-SGD, each of length $N=10^{7}$. The shaded lines depict individual runs of reg-SGD. The solid lines show the average errors for reg-SGD. The different colors correspond to various choices of $\alpha_{1}\in\{0.01,0.1,1,2\}$.
@@
We implemented both SGD and reg-SGD by partitioning the forward operator $A\in\mathbb{R}^{K\times d}$ into rows $A_{i}\in\mathbb{R}^{1\times d}$, $i=1,\dots,K$. Hence, $A_{i}x$ corresponds to the discretized ODE solution at location $s_{i}$. At each iteration, a batch of $16$ locations $(s_{i_{1}},\dots,s_{i_{16}})$ are sampled uniformly at random, and the gradient of $f$ is approximated by $\nabla f(x)\approx\frac{1}{16}\sum_{j=1}^{16}A_{i_{j}}^{\top}(A_{i_{j}}x-y_{i_{j}})\in\mathbb{R}^{d}$ and additionally perturbed by independent noise following a multivariate normal distribution with zero mean and covariance $0.001^{2}\cdot{\mathrm{Id}}$. For SGD we chose the step-size schedule $\alpha_{k}=100k^{-1/2}$. For reg-SGD we chose $\alpha_{k}=100k^{-2/3}$ and regularization $\lambda_{k}=0.001k^{-1/3}$. We initialize both algorithms for each repetition at zero.
In Figure˜12, we compare the expected and pathwise optimality gap (left) as well as the $L^{2}$ and pathwise error to the minimum-norm solution (right). While SGD shows fast convergence in terms of the optimality gap, it again fails to converge to the minimum-norm solution. In contrast, reg-SGD slows down the convergence in terms of the optimality gap, but safely reconstructs the minimum-norm solution. In Figure˜13 (left), we plot the reconstruction of the unknown right-hand side $x^{\dagger}$ resulting from the minimum-norm solution $x_{\ast}=A^{\dagger}y$, and from the last iterates of SGD and reg-SGD. Moreover, in Figure˜13 (right) we plot the corresponding ODE solutions when solving (11) with the estimated right-hand side.
@@ Appendix B Auxiliary results
Suppose that $f$ satisfies Section˜1.1, then the following statements hold true:
@@
Under Section˜1.1 the following estimates are satisfied for all $x\in\mathcal{X}$ and $\lambda\geq 0$:
@@
Let $C\geq 0$ be a constant such that for all $k\in\mathbb{N}$ one has $Y_{k}\geq-C$, almost surely. Set $(\tilde{Y}_{k})_{k\in\mathbb{N}}=(Y_{k}+C)_{k\in\mathbb{N}}$ and note that (13) still holds when replacing $(Y_{k})_{k\in\mathbb{N}}$ by $(\tilde{Y}_{k})_{k\in\mathbb{N}}$. Thus, the statement follows from Appendix˜B for the choice $\gamma_{k}\equiv 0$, $X_{k}\equiv 0$ and $(Z_{k})_{k\in\mathbb{N}}=(z_{k})_{k\in\mathbb{N}}$. ∎
Corollary B.7.
Let $(\mathcal{F}_{k})_{k\in\mathbb{N}}$ be a filtration and $(Y_{k})_{k\in\mathbb{N}}$, $(A_{k})_{k\in\mathbb{N}}$, $(B_{k})_{k\in\mathbb{N}}$ and $(C_{k})_{k\in\mathbb{N}}$ be non-negative and adapted processes satisfying almost surely that
$\sum_{k=1}^{\infty}A_{k}=\infty\quad\text{ , }\quad\sum_{k=1}^{\infty}B_{k}<\infty\quad\text{and}\quad\sum_{k=1}^{\infty}C_{k}<\infty\,.$
@@
Proof.
The proof follows the same lines as the proof of Lemma A.2 in [72]. For completeness, we provide the full details. Compared to Appendix˜B, we have $Y_{k}=Y_{k}$, $X_{k}=A_{k}Y_{k}$, $Z_{k}=B_{k}$ and $\gamma_{k}=C_{k}$. Using Appendix˜B we obtain the existence of $Y_{\infty}$ which is almost surely finite, integrable and satisfies $Y_{n}\to Y_{\infty}$ almost surely. Additionally, we have that $\sum_{k=1}^{\infty}X_{k}=\sum A_{k}Y_{k}<\infty$ implying that $\liminf_{k\to\infty}Y_{k}=0$, where we have used the assumption $\sum_{k=1}^{\infty}A_{k}=\infty$ almost surely. Since the limit inferior and limit coincide for converging sequences, the assertion follows by
$Y_{\infty}=\lim_{k\to\infty}Y_{k}=\liminf_{k\to\infty}Y_{k}=0\quad\text{almost surely}\,.$
Appendix C Finite-sum problems
In this section, we prove (6) from Section˜1.1 in the introduction. We consider the finite-sum optimization problem
$\min_{x\in\mathbb{R}^{d}}f(x)=\frac{1}{N}\sum_{i=1}^{N}f_{i}(x),$
where, for all $i=1,\dots,N$, $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $L_{i}$-smooth. The mini-batch estimator with mini-batch size $M\in\mathbb{N}$ is defined via $g_{k}=\frac{1}{M}\sum_{i\in M}\nabla f_{I_{i,k}}(X_{k-1}),$ for all $k\in\mathbb{N}$, where $(I_{i,k})_{i,k\in\mathbb{N}}$ is a family of iid. random variables that are uniformly distributed on $\{1,\dots,N\}$. The corresponding gradient noise is defined as $D_{k}=\frac{1}{M}\sum_{i\in M}(\nabla f_{I_{i,k}}(X_{k-1})-\nabla f(X_{k-1})).$ We show that in the finite-sum situation the ABC-condition, Section˜1.1, is satisfied. The following lemma is a version of [34, Lemma 4.20] with improved constants.
Lemma C.1.
The sequence $(D_{k})_{k\in\mathbb{N}}$ satisfies for all $k\in\mathbb{N}$
$\mathbb{E}\left[\|D_{k}\|^{2}\mid\mathcal{F}_{k-1}\right]\leq\frac{4L}{M}\bigl(f(X_{k-1})-f(x_{\ast})\bigr)+\frac{2\sigma_{\ast}^{2}}{M},$
where $\bar{L}=\frac{1}{N}\sum_{i=1}^{N}L_{i}$ and $\sigma_{\ast}^{2}=\frac{1}{N}\sum_{i=1}^{N}\|\nabla f_{i}(x_{\ast})\|^{2}$.
Proof.
Since, for all $i\in\{1,\dots,N\}$, $f_{i}$ is convex and $L_{i}$-smooth we get for all $x,y\in\mathbb{R}^{d}$ that
$f_{i}(x)-f_{i}(y)\leq\langle\nabla f_{i}(y),x-y\rangle+\frac{L_{i}}{2}\|x-y\|^{2}.$
For fixed $y\in\mathbb{R}^{d}$ let
$\varphi_{i}(x)=f_{i}(x)-f_{i}(y)-\langle\nabla f_{i}(y),x-y\rangle.$
Due to convexity of $f_{i}$, $\varphi_{i}$ is non-negative. Moreover, $\nabla\varphi_{i}(x)=\nabla f_{i}(x)-\nabla f_{i}(y)$ is $L_{i}$-Lipschitz. Thus, for $z=x-\frac{\nabla\varphi_{i}(x)}{L_{i}}$
$0$ $\leq\varphi_{i}(z)=\varphi_{i}(x)-\langle\nabla\varphi_{i}(x),\frac{\nabla\varphi_{i}(x)}{L_{i}}\rangle+\frac{L_{i}}{2}\|\frac{\nabla\varphi_{i}(x)}{L_{i}}\|^{2}$
$=f_{i}(x)-f_{i}(y)-\langle\nabla f_{i}(y),x-y\rangle-\frac{1}{2L_{i}}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2},$
which yields
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}\leq 2L_{i}(f_{i}(x)-f_{i}(y)-\langle\nabla f_{i}(y),x-y\rangle).$ (14)
Thus,
$\mathbb{E}\left[\|D_{k}\|^{2}\mid\mathcal{F}_{k-1}\right]$ $=\frac{1}{NM}\sum_{i=1}^{N}\|\nabla f_{i}(X_{k-1})-\nabla f(X_{k-1})\|^{2}$
$=\frac{1}{NM}\sum_{i=1}^{N}\|\nabla f_{i}(X_{k-1})-\nabla f_{i}(x_{\ast})-\nabla f(X_{k-1})+\nabla f_{i}(x_{\ast})\|^{2}$
$=\frac{2}{M}\sigma_{\ast}^{2}+\frac{2}{NM}\sum_{i=1}^{N}\|\nabla f_{i}(X_{k-1})-\nabla f_{i}(x_{\ast})-\nabla f(X_{k-1})\|^{2}$
Since $\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}(X_{k-1})-\nabla f_{i}(x_{\ast})=\nabla f(X_{k-1})$, we can use (14) with $x=X_{k-1}$ and $y=x_{\ast}$ to get
$\frac{1}{N}$ $\sum_{i=1}^{N}\|\nabla f_{i}(X_{k-1})-\nabla f_{i}(x_{\ast})-\nabla f(X_{k-1})\|^{2}\leq\frac{1}{N}\sum_{i=1}^{N}\|\nabla f_{i}(X_{k-1})-\nabla f_{i}(x_{\ast})\|^{2}$
$\leq 2\bar{L}f(X_{k-1})-f(x_{\ast})-\underbrace{\frac{1}{N}\sum_{i=1}^{N}\langle\nabla f_{i}(x_{\ast}),X_{k-1}-x_{\ast}\rangle}_{\langle\nabla f(x_{\ast}),X_{k-1}-x_{\ast}\rangle=0}.$
Appendix D Proofs of the main results
@@
Given Appendix˜B, this process $(E_{k})_{k\in\mathbb{N}_{0}}$ serves as Lyapunov function for computing the convergence rates stated in Section˜2.
Proposition D.1.
Suppose that Section˜1.1 and Section˜1.1 are fulfilled and and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with predictable (random) step-sizes and regularization parameters that are uniformly bounded from above and such that $(\lambda_{k})_{k\in\mathbb{N}}$ is almost surely decreasing. For $k\in\mathbb{N}$ denote by $\mathbb{A}_{k}=\{\alpha_{k}\leq\frac{2}{L+\lambda_{k}}\}\in\mathcal{F}_{k-1}$. Then, for all $k\in\mathbb{N}$,
$\mathbb{E}[1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Big(1-2\lambda_{k}\alpha_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big)1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}$
@@
Using Section˜1.1, the property $f\leq f_{\lambda_{k}}$, and the descent condition (12) applied to the $(L+\lambda_{k})$-smooth function $f_{\lambda_{k}}$, yields, for $k\in\mathbb{N}$,
$\mathbb{E}[1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}f_{\lambda_{k}}(X_{k})\mid\mathcal{F}_{k-1}]$ $\leq 1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}\Big(f_{\lambda_{k}}(X_{k-1})-\alpha_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)\|\nabla f_{\lambda_{k}}(X_{k-1})\|_{\mathcal{X}}^{2}$
@@
$\leq 1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}\Big(f_{\lambda_{k}}(X_{k-1})-\alpha_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)\|\nabla f_{\lambda_{k}}(X_{k-1})\|_{\mathcal{X}}^{2}$
@@
$\leq 1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}\Big(f_{\lambda_{k}}(X_{k-1})-\alpha_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)\|\nabla f_{\lambda_{k}}(X_{k-1})\|_{\mathcal{X}}^{2}$
@@
where in the last step we also used Appendix˜B (iii). Since each $f_{\lambda_{k}}$ is $\lambda_{k}$-strongly convex, it satisfies the Polyak-Łojasiewicz inequality
@@
$\mathbb{E}[1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}f_{\lambda_{k}}(X_{k})\mid\mathcal{F}_{k-1}]$ $\leq 1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}\Bigl(f_{\lambda_{k}}(X_{k-1})-2\alpha_{k}\lambda_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)(f_{\lambda_{k}}(X_{k-1})-f_{\lambda_{k}}(x_{\lambda_{k}}))$
@@
since $f_{\lambda_{k+1}}(X_{k})-f_{\lambda_{k}}(X_{k})\leq 0$. Combining the previous computations and using Appendix˜B (ii) yields
$\mathbb{E}[1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Bigl(1-2\lambda_{k}\alpha_{k}\Bigl(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Bigr)\Bigr)1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}+f_{\lambda_{k}}(x_{\lambda_{k}})-f_{\lambda_{k+1}}(x_{\lambda_{k+1}})$
$\quad+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}\bigl(1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}A(f_{\lambda_{k}}(X_{k-1})-f_{\lambda_{k}}(x_{\lambda_{k}}))+A\frac{\lambda_{k}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+C\bigr)$
$\leq\Bigl(1-2\lambda_{k}\alpha_{k}\Bigl(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Bigr)\Bigr)1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}+\frac{\lambda_{k}-\lambda_{k+1}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}$
$\quad+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}\Bigl(1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}AE_{k-1}+A\frac{\lambda_{k}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+C\Bigr).$
With the help of the energy function $(E_{k})_{k\in\mathbb{N}_{0}}$, we can bound the optimality gap of the true objective function, as well as the distance to the unique minimizer of the regularized objective function. For this, we rephrase Appendix˜B in the notation used in this section.
Lemma D.2.
Suppose that Section˜1.1 is fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4). Then the following estimates are satisfied for all $k\in\mathbb{N}$:
@@
    $\|X_{k}-x_{\lambda_{k+1}}\|_{\mathcal{X}}^{2}\leq\frac{2E_{k}}{\lambda_{k+1}}$.
D.1 General convergence result
First, we prove the general convergence results. We emphasize that no boundedness assumption is imposed on the reg-SGD scheme. In fact, in the proof of Theorem˜2.1 we will show that Assumptions 1.1-1.1 together with (17) imply that $\sup_{k\in\mathbb{N}_{0}}\|X_{k}\|_{\mathcal{X}}^{2}<\infty$ almost surely. The proof uses ideas from Theorem 4.1 in [56] and Lemma 3.1 in [28]. Due to the discretization error, we introduce and analyze a combined Lyapunov function $(\varphi_{k}+E_{k})_{k\in\mathbb{N}}$, where $\varphi_{k}=\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ and $E_{k}$ is defined in (15), in order to prove a descent step.
D.1.1 Proof of Theorem˜2.1